A Numerical Approach to Three-Dimensional Dendritic Solidification

Ryo Kobayashi

Abstract

We consider pattern formation during the supercooling solidification of a pure material, using a phase field model. The model gives rise to a rich variety of three-dimensional patterns, including very realistic dendritic crystal forms. We show how the strength of anisotropy has a crucial influence on the shape of crystals.